OpenStudy (anonymous):

For a particular sequence, T1 = 3and Tn = -2Tn-1 +1 then T5 equals? A: 7 B: 21 C: 11 D: 43

OpenStudy (anonymous):

@Chlorophyll Think you could help?

OpenStudy (anonymous):

Replace n = 5, can you?

OpenStudy (anonymous):

I have no idea.

OpenStudy (anonymous):

I'll rewrite the question so it looks clearer.

OpenStudy (anonymous):

Is it Tn-1 has both the subscript n-1?

OpenStudy (anonymous):

For a particular sequence, \[T _{1} = 3\] and \[T _{n} = -2t _{n-1}\] + 1, then \[T _{5}\] equals:

OpenStudy (anonymous):

It's just like that x_X

OpenStudy (anonymous):

I can't think of the shortcut !!! Here's the simple way: T2 = -2T1 +1 = -2 * 3 + 1 = -5 T3 = -2T2 +1 = -2 *(-5) +1 = 11 T4 = -2T3 +1 = -2 * 11+1 = -21 T5 = -2T4 +1 = -2*(-21)+1 = 42